OpenStudy (anonymous):

Can anyone calculate 1000000000 X(10 000^1000000000) in standard form? 0_0

OpenStudy (ash2326):

1000000000 if we count the no. of zeroes in this, it'll be 9 so this is 10^9 the other term is (10000^1000000000) 10000=10^4 1000000000= 10^9 so we have \[{10}^9*({(10^4)}^{10^9})\] now \[10^9*({10}^{4*10^9})\] so it can be simplified as \[{10}^{9+4*10^{9}}\]

OpenStudy (mathmate):

\[=10^{4000000009}\]